OpenStudy (anonymous):

Identify the horizontal asymptote of f(x) = (x^2+5x-3)/(4x-1)

OpenStudy (anonymous):

set the denominator equal to zero and solve i.e. solve for \(x\) : \(4x-1=0\)

OpenStudy (cwrw238):

won't it be vertical asymptote?

OpenStudy (amistre64):

horizontals are for large values of x

OpenStudy (amistre64):

for large values of x; this is practically the same as x^2/4x

OpenStudy (cwrw238):

vertical asymptote at x = 1/4?

OpenStudy (amistre64):

x/4 + 21/16 -------------- 4x-1 | x^2+5x-3 -(x^2-x/4) ----------- 21/4 x - 3 - (21/4x - 21/16) ----------------- R the limit as x appraoches infinity, the remainder goes to zero and we are left with ha = \(\frac{1}{16} (4x + 21)\) if im remembering it correctly

OpenStudy (amistre64):

err, thats the slant asymptote; there is non horizontal that i can tell